OpenStudy (rayman5766):

What is the solution to the equation? x – 1.2 = 3.6

OpenStudy (rayman5766):

wow u showed up nice i have one more

OpenStudy (texaschic101):

x - 1.2 = 3.6 --- add 1.2 to both sides x - 1.2 + 1.2 = 3.6 + 1.2 -- combine like terms x = 3.6 - 1.2 x = 2.4

OpenStudy (texaschic101):

oops...wait..I messed up

OpenStudy (texaschic101):

x = 3.6 + 1.2 x = 4.8

OpenStudy (texaschic101):

sorry about that....haven't had my coffee yet

OpenStudy (rayman5766):

oh ok i had that the first time

OpenStudy (texaschic101):

got another one ?

OpenStudy (rayman5766):

yea

OpenStudy (texaschic101):

lets see it

OpenStudy (rayman5766):

Use a transformation to solve the equation. t – (–16.5) = –13.4 A. t = –29.9 B. t = –3.1 C. t = 3.1 D. t = 29.9

OpenStudy (texaschic101):

I do not understand your question

OpenStudy (rayman5766):

u have to leav and join back it works

OpenStudy (texaschic101):

is it : t - 16.5 = 13.4 or is it t + 16.5 = 13.4 ???

OpenStudy (rayman5766):

ok do u see a bunch of squares if u do leave and come back

OpenStudy (rayman5766):

its 16.5 and 13.4

OpenStudy (texaschic101):

ok...hold on

OpenStudy (texaschic101):

t - (-16.5) = -13.4 --- distribute through the parenthesis t + 16.5 = -13.4 -- subtract 16.5 from both sides t + 16.5 - 16.5 = -13.4 - 16.5 -- combine like terms t = - 29.9

OpenStudy (rayman5766):

ok hold on lets see what i got 80% i passed thanks man

OpenStudy (texaschic101):

80 is not too bad...good for you :)

OpenStudy (texaschic101):

you deserve a medal for that :)

OpenStudy (rayman5766):

true i think i will do it agin later but u know any ways i have to go u can save some one els now

OpenStudy (texaschic101):

lol....I will try